OpenStudy (anonymous):

Americans are recycling about 61.9% of all yard trimmings. What amount of yard trimmings did the average American recycle per day in 2005? (Use the information in Exercise 14.) here is #14 In 2005, the average American generated 4.5 lb of waste per day. How much of that was paper and cardboard? yard trimmings 13.1% , paper and cardboard is 34.2%... my teacher says : first find 13.1% of the given total 4.5;then find 61.9% of that number for the answer but im not understanding it.

OpenStudy (anonymous):

The yard trimmings before recycling are part of the daily waste that an American produces

OpenStudy (anonymous):

So first we have to find how much of the total waste is equal to the yard trimmings.

OpenStudy (anonymous):

Do you see what part this is in the procedure that your teacher provided?

OpenStudy (anonymous):

i got a big graph?

OpenStudy (anonymous):

you shouldn't need a graph for this part.

OpenStudy (anonymous):

she wrote me back said first find 13.1% of the given total 4.5; then find 61.9% of that number for the answer .. i meant i got a chart

OpenStudy (anonymous):

So let's do those calculations

OpenStudy (anonymous):

The first part is simply 13.1%*4.5lb = 0.131 * 4.5 do you understand why we are doing this?

OpenStudy (anonymous):

not really sorry

OpenStudy (anonymous):

i will try to explain again. Of the total waste that a person produces, we know that 13.1 % of that waster is yard trimmings.

OpenStudy (anonymous):

The total amt of waste is 4.5 lb/day

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Therefore to calculate the amt of yard trimmings produced each day we have to take 13.1 % of the 4.5 lb/day

OpenStudy (anonymous):

does that make any sense?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

alright so this calculation 0.131*4.5 will give us the yard trimmings per day 0.131*4.5=0.5895 lbs of yard trimmings per day

OpenStudy (anonymous):

Following?

OpenStudy (anonymous):

yea cool

OpenStudy (anonymous):

alright, so this question asks us to find out how many yard trimmings are being recycled per day.

OpenStudy (anonymous):

We already know...

OpenStudy (anonymous):

Amount of yard trimmings produced each day = 0.5895 lb/day Percentage of yard trimmings that are recycled = 61.9%

OpenStudy (anonymous):

can you write the equation that will give us our final answer?

OpenStudy (anonymous):

how would i do that?

OpenStudy (anonymous):

by doing the 2nd thing that your teacher say to do.

OpenStudy (anonymous):

would i divide the two or ? thats what i have trouble with

OpenStudy (anonymous):

i don't know where this division by two is coming from. However all we are doing is taking 61.9% of the first number we found (0.5895 lb/day)

OpenStudy (anonymous):

so i subtract

OpenStudy (anonymous):

no, when we take % of a number we multiply

OpenStudy (anonymous):

50%*100 cents = 0.5*100 = 50 cents 25% of 1 hour = 25% * 60 minutes = 0.25* 60 minutes = 15 minutes

OpenStudy (anonymous):

so i 61.9% x05895?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

61.9%*0.05895

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

61.9%*0.5895

OpenStudy (anonymous):

36.49?

OpenStudy (anonymous):

not quite, before we multiply we need to change the percentage into a decimal. This is done by dividing the percentage by 100 75% = 75/100 = 0.75 25% = 25/100 = 0.25 30% =30/100 = 0.3

OpenStudy (anonymous):

so you would divide 61.9 by 100 and then multiply this number by 0.5895

OpenStudy (anonymous):

0.364

OpenStudy (anonymous):

yup, and this makes more sense since our answer should be smaller than 4.5

OpenStudy (anonymous):

and smaller than 0.5895

OpenStudy (anonymous):

smaller than 4.5 would be my answer?

OpenStudy (anonymous):

0.364 should be the answer, the units are lb/day

OpenStudy (anonymous):

wow cool thanks for taking your time to helping me understand it better and working with me

OpenStudy (anonymous):

of course, enjoy your day!

OpenStudy (anonymous):

yoou do the same